# The cosmic trimmer: Black-hole hair in scalar-Gauss-Bonnet gravity is altered by cosmology

### v1

@@ IV The good example: Cubic Galileon
A fair question is whether the type of problem appears in other scalar-tensor theories. This turns out not to be the case in general. As an example of a model which is free of this particular issue we now consider the Cubic Galileon model discussed in Ref. Babichev and Esposito-Farèse (2013), with action
@@
Notice that the kinetic term has a ghostly sign, which leads to self-accelerated cosmological solutions Deffayet et al. (2010). The homogeneous solution for the scalar of the form (21), asymptotes toward
@@
Now, turning on the mass of the black hole, we consider as previously the ansatz (10) for the Schwarzschild-de Sitter background metric. The analogue of (11) in this case is
$J_{r}=\varphi^{\prime}+\frac{1}{\Lambda^{2}}\left(\frac{Xf^{\prime}}{f}-\frac{2f\varphi^{\prime 2}}{r}\right)=\frac{C}{r^{2}f}\,.$ (26)
The above expression cannot be satisfied at all unless $X$ is finite, implying the solution of the form (13). This means that the value of $X$ is necessarily finite at a horizon for any $C$. This is in contrast to the case of sGB theory, where we had to fix $C$ to avoid the divergence of $X_{h}$. The value of $X_{h}$ at a horizon is found from (26) by taking the limit $r\to r_{h}$,
$\frac{C}{r_{h}^{2}}=\epsilon_{h}q-\frac{1}{\Lambda^{2}}\left(-X_{h}f_{h}^{\prime}+\frac{2q^{2}}{r_{h}}\right)+O(f)\,,$ (27)
Applying the same argument as below Eq. (22), also forces $q=q_{c}$ when homogeneity is required in the $r\to\infty$ limit. This is true for any $C$.
@@ V Extra operators alleviate the problem
The fact that the example of the Cubic Galileon works well, in contrast to the sGB case, can be traced to the form of the current (26). The explicit presence of $X$ in this equation forces $X_{h}$ to be finite. In this respect, the sGB operator is rather unique among the Horndeski operators in not providing an explicit $X$ term in the current (11). The natural question is then whether the presence of any of these additional operators may mitigate the problem faced by the vanilla sGB theory (1).
To have an explicit example, it is not difficult to check that when combining the sGB term with the Cubic Galileon term, the solution for the scalar field remains of the form (13), thus providing a finite $X_{h}$ at both horizons. We expect similar results when we combine sGB with other Horndeski terms that contain a function of $X$. Indeed, assuming the behaviour (13), for the sGB theory with an extra shift-symmetric term, the equation of motion for the scalar at a horizon becomes (cf. Eq. (27)),
$\frac{C}{r_{h}^{2}}=-\epsilon_{h}q+F\left(q,X_{h}\right)+\frac{4\alpha}{r_{h}^{3}}+O(f),$ (28)
where $F$ depends on the four Horndeski functions $G_{i}$. As one can see, for an arbitrary value of $C$, the solution $\varphi^{\prime}$ exists with finite $X$ at horizons and our modified sGB behaves just like the Cubic Galileon. It must be stressed that our argument does not establish for which particular values of $C$ solutions regular in the whole intermediate spacetime actually exist. Determining when this construction is viable is beyond the scope of this letter.
One may still think that picking $C$ as in the static case (7) would give physics similar to the static case. This is not true. Notwithstanding whether such a solution would be regular, the background for $\phi$ would now be time dependent, $F$ would have to make a significant contribution to solve the two-horizon problem and, as a consequence, the perturbations would have very different dynamics. Observables such as radiated power and the quasi-normal-mode spectrum would be strongly affected.
@@ VI Discussion & Conclusions
Let us take a different look at what has happened here: it is instructive to discuss shift-symmetric theories like sGB in terms of the dynamics of the corresponding shift charges – the equation of motion for the scalar (3) is a conservation equation for the shift-charge current, $J^{\mu}$ Pujolas et al. (2011), with $J^{0}$ the shift-charge density and $J^{i}$ is the spatial current giving the motion of these charges. The static sGB solution then corresponds to a static radial current (5) with flux $C$ across any surface enclosing the horizon. [^2] We stress that, in the shift-charge picture, $C$ is a flux of conserved shift charges the density of which is given by $J^{0}$, not a scalar charge as it is usually referred to. For this static configuration to be maintained, the charges must leave the spacetime at the rate they enter, and they are able to do so at spatial infinity. Keeping $X$ finite on a horizon means that the horizon must produce or absorb a particular flux $C$ of charges given by (7). Introducing a second horizon, e.g. cosmology, gives a different flux requirement. The two are in general incompatible, so the charges will begin to amass in the intermediate spacetime between the horizons breaking the staticity assumption of the solution.
@@
Cosmology introduces time dependence to the scalar, with an attractor solution $q=q_{c}$ (20). This is also a vacuum solution with $J^{\mu}=0$ to which any black hole would have to match. This requires stationary time-dependent hair (16). However, the requirement of finiteness on both horizons requires a $q$ grossly different from $q_{c}$ and therefore a non-zero shift-charge density. This cannot be matched to a homogeneous cosmology.
@@
We have not found any stationary solutions in sGB under our assumptions, but it may well be that the evolution away from it is slow and therefore one might be able to use the non-evolving $q=q_{c}$ solution with $C$ given by (16) for sufficiently long to study e.g. binary inspirals. Even so the static solution is not a good approximation since the cosmological background provides a medium which will affect the dynamics.
@@
Alternatively, the introduction of (an) extra Horndeski terms resolves the inconsistency of the requirements put on the solution. The presence of $X$ in Eq. (28) means the theory is less rigid and allows for any choice of flux $C$ across any horizon even in the presence of the $q_{c}$ compatible with the cosmological vacuum needed to be reached at large distances. This is of course not sufficient that guarantee that such a solution exists – it only provides the boundary conditions on the two horizons which must then be connected through a non-singular source-free solution. This underlines how spectacularly the static sGB solution in cosmology fails – it cannot even support the required boundary conditions.
@@
The price of this well-behaved alternative is that the hair which is necessary once a horizon appears in sGB Eq. (7) is no longer the same upon the introduction of any generic additional Horndeski operator. In any such good solution, the background that this hair gives for the fluctuations will be drastically different to the static sGB solution and we expect will strongly affect the predictions of any physical process. We leave the construction of viable configuration and the study of its properties to a separate work.
@@ VII Acknowledgments
The authors would like to thank Christos Charmousis, Gilles Esposito-Farèse and Nicolas Lecoeur for helpful discussions. E.B. acknowledges the support of ANR grant StronG (ANR-22-CE31-0015-01). The work of L.G.T. was supported by the European Union (Grant No. 101063210). I.S. is supported by the Czech Science Foundation (GAČR) project PreCOG (Grant No. 24-10780S). E.B. and L.G.T. would like to thank the hospitality of the Centro de Ciencias de Benasque Pedro Pascual during the initial stages of this work.

### v2

@@ IV The good example: Cubic Galileon
A fair question is whether this type of problem appears in other scalar-tensor theories. This turns out not to be the case in general. As an example of a model which is free of this particular issue, we now consider the cubic Galileon model discussed in Ref. [24], with the action
@@
Notice that the kinetic term has a ghostly sign, which leads to self-accelerated cosmological solutions [25]. The homogeneous solution for the scalar of the form (20) asymptotes toward
@@
Now, turning on the mass of the black hole, we consider as previously the ansatz (9) for the Schwarzschild-de Sitter background metric. The analogue of Eq. (10) in this case is
$J_{r}=\varphi^{\prime}+\frac{1}{\Lambda^{2}}\left(\frac{Xh^{\prime}}{h}-\frac{2f\varphi^{\prime 2}}{r}\right)=\frac{C}{r^{2}\sqrt{hf}}\,.$ (25)
The above expression cannot be satisfied at all unless $X$ is finite, implying a solution of the form (12). This means that the value of $X$ is necessarily finite at a horizon for any $C$. This is in contrast to the case of sGB theory, where we have to fix $C$ and $q$ to avoid the divergence of $X_{h}$. The value of $X_{h}$ at a horizon is found from Eq. (25) by taking the limit $r\to r_{h}$,
$\frac{C}{r_{h}^{2}}=\epsilon_{h}q+\frac{1}{\Lambda^{2}}\sqrt{\frac{f}{h}}\left(X_{h}h^{\prime}-\frac{2q^{2}}{r_{h}}\right)+O(h,f)\,.$ (26)
This imposes no condition on either $C$, $q$, or $X_{h}$, which may or may not be subject to other constraints away from the horizons.
@@ V Extra operators alleviate the problem
The fact that the example of the cubic Galileon works well, in contrast to the sGB case, can be traced to the form of the current (25). The explicit presence of $X$ in this equation forces $X_{h}$ to be finite. In this respect, the sGB operator is rather unique among the Horndeski operators in not providing an explicit $X$ term in the current (10). The natural question is then whether the presence of any of these additional operators may mitigate the problem faced by the vanilla sGB theory [Eq. (1)].
To have an explicit example, it is not difficult to check that when combining the sGB term with the cubic Galileon term, the solution for the scalar field remains of the form (12), thus providing a finite $X_{h}$ at both horizons. We expect similar results when we combine sGB with other Horndeski terms that contain a function of $X$. Indeed, assuming the behavior in Eq. (12), for the sGB theory with an extra shift-symmetric term, the equation of motion for the scalar at a horizon becomes [cf. Eq. (26)],
$\frac{C}{r_{h}^{2}}=-\epsilon_{h}q+F\left(q,X_{h}\right)+4\alpha\sqrt{\frac{f}{h}}\frac{h^{\prime}}{r_{h}^{2}}+O(h,f),$ (27)
where $F$ depends on the four Horndeski functions $G_{i}$. As one can see, for arbitrary values of $C$ and $q$, the solution $\varphi^{\prime}$ exists with finite $X$ at horizons and our modified sGB behaves just like the cubic Galileon. It must be stressed that our argument does not establish for which particular values of $C$ solutions regular in the whole intermediate spacetime actually exist. Determining when this construction is viable is beyond the scope of this letter.
One may still think that picking $C$ as in the static case (7) and $q=0$ would give physics similar to the static case. This is not true. Notwithstanding whether such a solution would be regular, the background for $\phi$ would now be very different, as $F$ would have to make a significant contribution to solve the two-horizon problem and, as a consequence, the perturbations would have very different dynamics. GW observables such as radiated power and the quasi-normal-mode spectrum would be strongly affected.
@@ VI Discussion & Conclusions
Let us take a different look at what has happened here: it is instructive to discuss shift-symmetric theories like sGB in terms of the dynamics of the corresponding shift charges – the equation of motion for the scalar (3) is a conservation equation for the shift-charge current, $J^{\mu}$ [26], where $J^{0}$ is the shift-charge density and $J^{i}$ is the spatial current giving the motion of these charges. The static sGB solution then corresponds to a static radial current (5) with flux $C$ across any surface enclosing the horizon. [^3] We stress that, in the shift-charge picture, $C$ is a flux of conserved shift charges the density of which is given by $J^{0}$, not a scalar charge as it is usually referred to. For this static configuration to be maintained, the charges must leave the spacetime at the rate they enter, and they are able to do so at spatial infinity. Keeping $X$ finite on a horizon means that the horizon must produce or absorb a particular flux $C$ of charges given by Eq. (7). Introducing a second horizon, e.g., cosmology, gives a different flux requirement. The two are in general incompatible, so the charges will begin to amass in the intermediate spacetime between the horizons breaking the staticity assumption of the solution.
@@
Cosmology introduces time dependence to the scalar, with an attractor solution $q=q_{c}$ [Eq. (19)]. This is also a vacuum solution with $J^{\mu}=0$ to which any black hole would have to eventually match. This requires stationary time-dependent hair [Eq. (15)]. However, the requirement of finiteness on both horizons requires a $q$ grossly different from $q_{c}$, and therefore a shift-charge cloud extending as far as $r\sim 1/H^{2}r_{s}$, much beyond the cosmological horizon.
@@
We have not found any reasonable stationary solutions in sGB under our assumptions, but it may well be that the evolution away from the static solution is slow, and therefore one might be able to use the $q\simeq q_{c}$ solution, with $C$ given by (15) for sufficiently long to study, e.g., binary inspirals. Even so, the static solution is not a good approximation, since the cosmological background provides a medium which may affect the dynamics.
@@
Alternatively, the introduction of (an) extra Horndeski terms resolves the inconsistency of the requirements put on the solution. The presence of $X$ in Eq. (27) means the theory is less rigid and allows for any choice of flux $C$ across any horizon, even in the presence of the $q\simeq q_{c}$ compatible with the cosmological vacuum needed to be reached at horizon scales. This is, of course, not sufficient to guarantee that such a solution exists – it only provides the boundary conditions on the two horizons, which must then be connected through a non-singular source-free solution. This underlines how spectacularly the static sGB solution in cosmology fails – it cannot even support the required boundary conditions.
@@
The price of this well-behaved alternative is that the hair which is necessary once a horizon appears in sGB Eq. (7) is no longer the same upon the introduction of any generic additional Horndeski operator. In any such good solution, the background that this hair gives for the fluctuations will be drastically different from the static sGB solution, and we expect that it will strongly affect the predictions of any physical process. In particular, predictions for GW signatures would need to be reassessed in light of our findings. We leave the construction of a viable configuration and the study of its properties to a separate work.
@@ VII Acknowledgments
The authors would like to thank Christos Charmousis, Gilles Esposito-Farèse, Dražen Glavan, and Nicolas Lecoeur for helpful discussions. E.B. acknowledges the support of ANR grant StronG (ANR-22-CE31-0015-01). The work of L.G.T. was supported by the European Union (Grant No. 101063210). I.S. is supported by the Czech Science Foundation (GAČR) project PreCOG (Grant No. 24-10780S). E.B. and L.G.T. would like to thank the hospitality of the Centro de Ciencias de Benasque Pedro Pascual during the initial stages of this work.